A box company produce 2 sizes of pizza boxes. The smaller box has a surface area of 260 square inches. The larger box has a length, width, and height that are twice those of the smaller box. What is the surface area of the larger box?
Answer by ankor@dixie-net.com(22740)

A box company produce 2 sizes of pizza boxes. The smaller box has a surface
area of 260 square inches. The larger box has a length, width, and height
that are twice those of the smaller box. What is the surface area of the larger box?
:
The surface area of the small rectangular box
2LW + 2LH + 2WH = 260
:
Simplify divide by 2
LW + LH + WH = 130
:
Big Box S.A, (doubling length, width, and height)
SA = 2(2L*2W) + 2(2L*2H) + 2(2W*2H)
SA = 8LW + 8LH + 8WH
:
factor out 8
SA = 8(LW + LH + WH)
:
from the small box we have: LW + LH + WH = 130, so we have
SA = 8(130)
SA = 1040 sq/in surface area of the big box